OpenStudy (anonymous):

Solve the equation. |k + 9| = 4 A. {–13, –5} B. {–5, 5} C. {–5, 13} D. {all real numbers greater than or equal to –13 and less than or equal to –5}

OpenStudy (mathstudent55):

To solve this type of an absolute value equation, just change it into this: k + 9 = 4 or k + 9 = -4 Now solve each simple equation. Then leave the two answers separated by the word "or."

OpenStudy (anonymous):

I'll fan and medal :)

OpenStudy (anonymous):

So is it B?

OpenStudy (amoodarya):

mathstudent55 give the right answer you just simplify them to ...

OpenStudy (mathstudent55):

No. Did you solve each equation? Let's do it together. k + 9 = 4 or k + 9 = -4 In each equation, you see that 9 is being added to k on the left side. To get rid of that 9, we must subtract 9 from both sides in both equations. k + 9 = 4 or k + 9 = -4 - 9 -9 - 9 -9 -------------------------- k = -5 or k = -13

OpenStudy (anonymous):

Oh, oh oh oh oh!! I get it so its C!?

OpenStudy (mathstudent55):

In the left equation, you get 4 - 9 = -5. In the right equation, you get -4 - 9 = -13. C. has POSITIVE 13. We got NEGATIVE 13 and -5. It's not C.

OpenStudy (anonymous):

Oh. wow. Sorry. It's A...

OpenStudy (mathstudent55):

Great. Now you got it.